Mole fraction of glucose is 0.25.Calculate molarity and molality if density is 1.5g/ml

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
From the Molarity formulae:
Molality= moles of solutex100/mass of dissolvable. 
mass of dissolvable =0.75x18( since dissolvable is water) = 13.5; 
m= 0.25*100/13.5=1.85m. 
Volume of water= 18/1.5 = 12ml. 
along these lines M= 0.25*100/12=2.083 moles/ml.